If 75% of any increase in income is spent on consumption, then an increase in autonomous investment of R1 billion results in an increase in national income of as much as


R5 billion


R1.33 billion


R4 billion


R6 billion


1
Expert's answer
2021-09-16T08:19:38-0400

MPC=0.75 or 75%=m

MPC= mC/mY

MPS+MPC=1

Y=C+I+G

but C=mY

Therefore, Y=mY+I+G

Factoring out; Y=(I+G)/(1-m)

Differentiating with respect to Investment:

therefore, Y=1/(1-m)

Y=1/0.25=4

this is the multiplier effect meaning that a unit increase in autonomous investment leads to an increase in national income by four units. therefore an increase in autonomous investment by R 1billion leads to increase in national income by R 4billion.